OpenStudy (anonymous):

The ratio of the radii of two spheres is 3:1. The volume of the larger sphere is 729in^3. What is the volume of the smaller sphere?

OpenStudy (anonymous):

volumes are again \[length^3\] so \[(3/1)^3 = 729/x\] \[27 = 729/x\] \[x = 27inches^3\]